Andrew drove at an average speed of 56 mph over six hours. During the first three hours, he drove at a speed that was sixty-eight mph less than two times the speed he averaged for the last three hours. How fast did he drive during the first three hours?
Answer by checkley77(12844)

3(2X-68)+3X=56*6
6X-204+3X=336
9X=336+204
9X=540
X=540/9
X=60 MPH. FOR THE SECOND PART OF THE TRIP.
2*60-68=120-68=52 MPH. FOR THE FIRST PART OF THE TRIP.
PROOF
3*52+3*60=56*6
156+180=336
336=336